# Capsule Network Projectors are Equivariant and Invariant Learners

### v3

@@ 5.2 Capsule Networks
CapsNets present an alternative architecture to CNNs, addressing their limitations by explicitly preserving hierarchical spatial relationships between features (Sabour et al., 2017). CapsNets replace scalar neurons with vector or matrix poses, representing specific concepts at different levels of a parse tree as the network goes deeper. The first layer (primary capsules) corresponds to the most basic parts, while capsules in deeper layers represent more complex concepts made up of the simpler concepts as they get closer to the final layer where each capsule corresponds to a specific class.
@@
The key novelty in CapsNets is the routing mechanism, which determines the contributions of lower-level capsules to higher-level capsules. Numerous routing algorithms, both iterative and non-iterative, have been proposed to address the efficiency and effectiveness of this process (Sabour et al., 2017; Mazzia et al., 2021; Feng et al., 2024; Ribeiro et al., 2020; Hinton et al., 2018; De Sousa Ribeiro et al., 2020; Yang et al., 2021; Everett et al., 2024). Among these, SRCaps (Hahn et al., 2019) introduces a non-iterative routing mechanism. This method maintains all of the desirable properties of CapsNets, such as equivariance while largely mitigating the time cost of iterative methods, at the price of a small amount of performance. However, SRCaps faces the same limitations as other CapsNets where high resolution or high class datasets are beyond the network’s abilities when trained in a standard fashion. For a more detailed description of capsule routing mechanisms please check this review (Ribeiro et al., 2022).
@@ 6 Conclusion
Our proposed method demonstrates how self-supervised CapsNets can be employed to better learn equivariant representations, leveraging architectural assumptions removing the the need to explicitly split representation vectors and train separate projector networks. The resulting solution, CapsIE, achieves state-of-the-art performance in equivariant downstream benchmarks with a improvement of 0.01 $R^{2}$ on prior self-supervised rotation prediction tasks. Our results contribute significantly to the avocation of CapsNets in self-supervised representation learning, introducing desirable properties with improved effectiveness over MLP projectors.
Ethics Statement. This work aims to learn higher quality and more applicable representations of images without human generated annotations, therefore such methods can lead to positive societal impacts the development of more accurate or informative models for a number of downstream tasks. However, as is the case with all vision systems, there is potential for exploitation and security concerns and one should take into consideration AI misuse when extending our method.
In addition, the use of CapsNets has shown little improvement when scaled, similar to the findings of (Everett et al., 2023; Mitterreiter et al., 2023; Nair et al., 2021) where the addition of more capsule layers has either led to stagnated or decreased performance. We strongly feel that our work should be revisited when a capsule network that can both scale and retain the desirable properties of capsule networks is found.
As stated in the original dataset proposal and the problem setting the methodology presented rely on the group elements being known. Hence, the applicability of the proposed method is only possible in settings where group elements are known. However, we present findings where additional group elements are preserved without prior knowledge of the group, this suggests the CapsNets are more capable than previous invariant methods at capturing equivariant properties, thus opening an intriguing direction for future work.
Reproducibility Statement. We have included the source code required to reproduce all experiments and ablation studies presented in the main body of the paper and the appendix. We also provide the details of our training protocol in the appendix.
Appendix A 3DIEBench Dataset
@@
Typical equivariant datasets are generally handcrafted and simple, with a small amount of classes and instances within each class. This is due to the time needed in order to ensure correctness. While standard image datasets do allow for testing invariance in the form of augmenting the same image in two different ways, they do not allow for precise transformation of the subject. Thus the need for a new, synthetic dataset.
We use the 3DIEBench (Garrido et al., 2023) dataset [^1] The full dataset and splits employed can be found at https://github.com/facebookresearch/SIE, which has been created specifically to be a hard yet controlled test-bed for invariant and equivariant methods. The dataset consists of 52,472 3d objects across 55 classes of 3d objects from ShapeNetCorev2 (Chang et al., 2015) posed in 50 different views as well as the latent information of the view, this can be seen in figure 5. For training we then randomly select two views from each model in the training set. The parameters by which the model could have been augmented can be seen in table 4.
Table 4: Values of the factors of variation used for the generation of 3DIEBench. Each value is sampled uniformly from the given interval. Object rotation is generated as Tayt-Bryan angles using extrinsic rotations. Light position is expressed in spherical coordinates. This table is sourced from (Garrido et al., 2023).
@@
Object rotation X $-\frac{\pi}{2}$ $\frac{\pi}{2}$
Object rotation Y $-\frac{\pi}{2}$ $\frac{\pi}{2}$
Object rotation Z $-\frac{\pi}{2}$ $\frac{\pi}{2}$
@@
Light $\theta$ 00 $\frac{\pi}{4}$
Light $\phi$ 00 $2\pi$
Appendix B Training protocols
B.1 CapsIE Pre-training
@@
For ablation studies, where we explicitly state, we train for fewer epochs and with a smaller batch size, 500 and 512 respectively. We find in practice this setting is a strong proxy for full training performance and significantly save computational resource.
Training time for 2000 epochs with batch size of 1024, as previously stated, took approximately 22 hours using 3 Nvidia A100 80GB GPU’s, with 64 capsule models taking approximately 25 hours using 6 Nvidia A100 80GB GPU’s.
B.2 Downstream Evaluation
In our work we perform evaluation on both the frozen ResNet-18 representations and the representations from the primary capsules layer, which are evaluated using either a linear classifier or additional capsule layer acting as a class capsules, i.e the number of capsules is set to the number of classes and activations are used as the logits. Here, we detail the exact training protocols to ensure complete reproducability.
@@ B.2 Downstream Evaluation
B.3 Supervised Training of SR-Caps
@@
Table 6: Training settings for our supervised Self Routing Capsule Network model. Settings are the same for all number of capsules. - denotes that this element is not used.
@@
B.4 Invariant and Equivariant SSL Benchmarks

### v4

@@ 5.2 Capsule Networks
CapsNets present an alternative architecture to CNNs, addressing their limitations by explicitly preserving hierarchical spatial relationships between features (Sabour et al., 2017). CapsNets replace scalar neurons with vector or matrix poses, representing specific concepts at different levels of a parse tree as the network goes deeper. The first layer (primary capsules) corresponds to the most basic parts, while capsules in deeper layers represent more complex concepts composed of simpler concepts as they get closer to the final layer, where each capsule corresponds to a specific class.
@@
The key novelty in CapsNets is the routing mechanism, which determines the contributions of lower-level capsules to higher-level capsules. Numerous routing algorithms, both iterative and non-iterative, have been proposed to address the efficiency and effectiveness of this process (Sabour et al., 2017; Mazzia et al., 2021; Feng et al., 2024; Ribeiro et al., 2020; Hinton et al., 2018; De Sousa Ribeiro et al., 2020; Yang et al., 2021; Everett et al., 2024). Among these, SRCaps Hahn et al. (2019) introduces a non-iterative routing mechanism. This method retains all the desirable properties of CapsNets, such as equivariance, while largely mitigating the time cost of iterative methods at the expense of a slight performance loss. However, SRCaps faces the same limitations as other CapsNets where high-resolution or high-class datasets are beyond the network’s abilities when trained in a standard fashion. For a more detailed description of capsule routing mechanisms, please check this review De Sousa Ribeiro et al. (2024).
5.3 Part-Whole Decomposition and Interpretability
Capsule networks naturally lend themselves to an explicit part–whole decomposition because each capsule encodes an entity’s instantiation parameters (pose, presence and related attributes) and lower-level capsules cast “votes” for higher-level capsules so that coherent clusters of votes identify which parts belong to which whole – an idea formalised as routing-by-agreement in the original capsule proposals (Sabour et al., 2017). This voting/routing mechanism produces structured, disentangled representations (pose matrices or vectors plus activation lengths) that are directly interpretable: pose parameters report geometric transformations while activation magnitudes reflect part/object presence, enabling inspection of what parts drove a decision (Hinton et al., 2018; De Sousa Ribeiro et al., 2024). Work on stacked/autoencoding capsule models has further shown that object-level capsules can be learned (even unsupervised) from part poses, strengthening the claim that capsules implement an intrinsic compositional (part→object) inductive bias (Kosiorek et al., 2019). Other formulations that treat routing as posterior inference or introduce explicit routing uncertainty demonstrate how probabilistic/variational routing makes the part–whole assignments and the model’s confidence in them explicit, improving robustness and giving a principled way to read out uncertainty about which parts compose a given object (De Sousa Ribeiro et al., 2020).
@@ 6 Conclusion
Our proposed method demonstrates how self-supervised CapsNets can be employed to better learn equivariant representations, leveraging architectural assumptions, removing the need to explicitly split representation vectors and train separate projector networks. The resulting solution, CapsIE, achieves state-of-the-art performance in equivariant downstream benchmarks with an improvement of 0.01 $R^{2}$ on prior self-supervised rotation prediction tasks. Our results contribute significantly to the application of CapsNets in self-supervised representation learning, introducing desirable properties with improved effectiveness over MLP projectors.
This work aims to learn higher quality and more applicable representations of images without human-generated annotations; therefore, such methods can lead to positive societal impacts and the development of more accurate or informative models for a number of downstream tasks. However, as is the case with all vision systems, there is potential for exploitation and security concerns; therefore, one should consider AI misuse when extending our method.
As stated in the original dataset proposal and the problem setting, the methodology presented relies on the group elements being known. Hence, the applicability of the proposed method is only possible in settings where group elements are known. We additionally explore alternative equivariant tasks that were not explicitly optimised for as a diagnostic task to evaluate if the capsule projector additionally encodes additional group elements. Our findings determine that our method retains the fine-grained control of learning equivariant embeddings specified by the objective function, a beneficial property presented in prior methods to avoid learning of arbitrary group elements.
While our capsule networks perform well on the equivariant tasks, the invariant performance has suffered as a consequence. We attribute this to the reduced capacity of capsule networks, where the number of embeddings is significantly lower than that of the MLP counterpart. To improve performance, alternative capsule networks other than SRCaps could be employed, which have shown improved classification performance, or the number of capsules can be scaled up, where Figure 4 demonstrated improved top-1 performance with greater capsule number. Lastly, a further limitation of our work is the absence of multiple seed runs in most experiments (with the exception of Table 4), consistent with (Bardes et al., 2022) and (Garrido et al., 2023), due to the disproportionately high computational costs.
Appendix A Appendix
A.1 3DIEBench Dataset
@@ Appendix A Appendix
Typical equivariant datasets are generally handcrafted and simple, with a small number of classes and instances within each class. This is due to the time needed in order to ensure correctness. While standard image datasets do allow for testing invariance in the form of augmenting the same image in two different ways, they do not allow for the precise transformation of the subject. Thus, there is a need for a new, synthetic dataset.
We use the 3DIEBench (Garrido et al., 2023) dataset [^2] The full dataset and splits employed can be found at https://github.com/facebookresearch/SIE, which has been created specifically to be a hard yet controlled test-bed for invariant and equivariant methods. The dataset consists of 52,472 3d objects across 55 classes of 3d objects from ShapeNetCorev2 (Chang et al., 2015) posed in 50 different views as well as the latent information of the view. This can be seen in figure 5. For training we then randomly select two views from each model in the training set. The parameters by which the model could have been augmented are listed in Table 6.
Table 6: Values of the factors of variation used for the generation of 3DIEBench. Each value is sampled uniformly from the given interval. Object rotation is generated as Tait-Bryan angles using extrinsic rotations. Light position is expressed in spherical coordinates. This table is sourced from (Garrido et al., 2023).
@@
Object rotation X $\frac{\pi}{2}$ $\frac{\pi}{2}$
Object rotation Y $\frac{\pi}{2}$ $\frac{\pi}{2}$
Object rotation Z $\frac{\pi}{2}$ $\frac{\pi}{2}$
@@
Light $\theta$ 0 $\frac{\pi}{4}$
Light $\phi$ 0 $2\pi$
A.2 3DIEBench Problem Statement
First, we define equivariance by defining a Group consisting of a set $G$ and a binary operation $\cdot$ on $G$, $\cdot:G\times G\rightarrow G$ such that $\cdot$ are associative; there is an identity $e$ which satisfies $e\cdot a=a=a\cdot e,\forall a\in G$; and for each $a\in G$ there exists an inverse $a^{-1}$ such that $a\cdot a^{-1}=e=a^{-1}\cdot a$. Group actions are concerned with how groups manipulate sets, where the left group action can be defined as a function $\alpha$ of group $G$ and set $S$, $\alpha:G\times S\rightarrow S$ such that $\alpha(e,s)=s,\forall s\in S$, and $\alpha(g,\alpha(h,s))=\alpha(gh,s),\forall s\in S,\text{and }\forall g,h\in G$. In our setting, we are concerned with group representations which are linear group actions acting on vector space $V$, which we define as $\rho:G\rightarrow GL(V)$ where $GL(V)$ is the general linear group on $V$. Here $\rho(g)$ describes the transformation applied to both the input data $x$ and latent $f(x)$ given parameters $g$ (Park et al., 2022; Garrido et al., 2023). Transformations comprise colour scaling and shifting, and rotations around a fixed point.
$f(\rho_{X}(g)\cdot x)=\rho_{Y}(g)\cdot f(x).$ (8)
The goal is to therefore learn $f$ and $\rho_{Y}$ to construct representations that are equivariant to viewpoint transformations when $\rho_{X}$ is not known, but the group elements $g$ that parameterise the transformations are known.
A.3 Training protocols
A.3.1 CapsIE Pre-training
@@
For ablation studies, where we explicitly state, we train for fewer epochs and with a smaller batch size, 500 and 512, respectively. We find in practice that this setting is a strong proxy for full training performance and significantly saves computational resources.
Training time for 2000 epochs with batch size of 1024, as previously stated, took approximately 22 hours using three Nvidia A100 80GB GPUs, with 64 capsule models taking approximately 25 hours using six Nvidia A100 80GB GPUs.
A.3.2 Downstream Evaluation
In our work, we perform evaluation on both the frozen ResNet-18 representations and the representations from the primary capsules layer, which are evaluated using either a linear classifier or an additional capsule layer acting as class capsules, i.e the number of capsules is set to the number of classes and activations are used as the logits. Here, we detail the exact training protocols to ensure complete reproducability.
@@ A.3.2 Downstream Evaluation
A.3.3 Supervised Training of SR-Caps
@@
Table 8: Training settings for our supervised Self Routing Capsule Network model. Settings are the same for all numbers of capsules. - denotes that this element is not used.
@@
A.3.4 Invariant and Equivariant SSL Benchmarks